OpenStudy (anonymous):

Hello- exponential equations! Have answers, need help with understanding! Thanks in advance. Solve for x. 1) 3^(2x+1)*5(2x-1) = 42 2) 18^(2x+1) = 5*6^(2x+1)

OpenStudy (anonymous):

ok so you can put Ln on both sides of the equation.. lets take equation number 2 for example : Ln(18^(2x+1)) = Ln(5*6^(2x+1)) do you think you can continue ?

OpenStudy (anonymous):

i'll try for a bit! just a sec (2x+1)ln18 = (2x+1)ln(5*6) is that right?

OpenStudy (anonymous):

so far i mean

OpenStudy (anonymous):

not quite .. your left side is right but look.. your right side should be : Ln(5*6^(2x+1)) = Ln(5) + (2x+1)Ln(6) Do you understand why ?

OpenStudy (anonymous):

i see, 5*6^2x+1 .... (take ln) = ln5*ln6^2x+1 = ln5 + (2x+1)ln6

OpenStudy (anonymous):

that already helps

OpenStudy (anonymous):

ok, i'll try myself for a bit

OpenStudy (anonymous):

wait .. writing : ln5*ln6^2x+1 isnt correct see Ln(a*b) = Ln(a) + Ln(b)

OpenStudy (anonymous):

ok keep going let me know if you need anything else

OpenStudy (anonymous):

thanks so far

OpenStudy (anonymous):

logs/ln follow distributive law with the coefficient right? (2x+1)ln6 can = (2x)ln6 + ln6

OpenStudy (anonymous):

sure

OpenStudy (phi):

ln(6) is just a number. Just like 2 only fancier (more digits)

OpenStudy (anonymous):

im new to logarithms so its easy to forget that :-) it looks fancy

OpenStudy (anonymous):

awesome, those tips helped, i get it now. i'll try the first one

OpenStudy (phi):

Here is one way to solve the 2nd one: (2x+1)*ln(18)= ln(5)+(2x+1)*ln(6) subtract (2x+1)*ln(6) from both sides: (2x+1)*ln(18) - (2x+1)*ln(6) = ln(5) factor out (2x+1) on the left side: (2x+1)*(ln(18)-ln(6))= ln(5) and though not necessary, notice ln(18)- ln(6)= ln(18/6)= ln(3), so 2x+1= ln(5)/ln(3) and x is x= (ln(5)/ln(3) -1)* 0.5

OpenStudy (anonymous):

that's exactly what i did too, cool but was making the mistake of applying ln improperly as coolsector pointed out, then also forgetting that lnx is "just a number" was making same mistakes with the first so it should go better

OpenStudy (anonymous):

in being natural log

OpenStudy (anonymous):

got both of them now! thank you both!